Global insurance has found that 20% (1 in 5) of all insurance policies are surrendered (cashed in) before their maturity date. Assume that 10 policies are randomly selected from the policies data base


What is the probability that:


Atleast 2 out of the 10 randomly selected policies will be surrendered before their maturity date?

1
Expert's answer
2022-03-01T07:17:55-0500

Let "X=" the numbers of insurance policies surrendered (cashed in) before their maturity date: "X\sim Bin (n, p)."

Given "n=10, p=0.2,q=0.8"


"P(X\ge2)=1-P(X=0)-P(X=1)"

"=1-\dbinom{10}{0}(0.2)^0(0.8)^{10-0}-\dbinom{10}{1}(0.2)^1(0.8)^{10-1}"

"=1-0.1073741824-0.268435456"

"=0.6241903616"